the period of |sinx|+|cosx| is

the period of
|sinx|+|cosx| is

Grade:12

4 Answers

Ritika Kapoor
13 years ago

the ans. is (pi)/2

since here we get f(p/2 +x) =f(x)

|sinx| bcums |cosx| n vice versa..

Mir Noaman Ali IIT-Roorkee
13 years ago

THE PERIOD OF MOD SIN X + MOD COS X IS PI/2 AS SIN X AND COS X ARE COMPLIMENTARY PAIRWISE COMPARIBLE FUNCTIONS

SAGAR SINGH - IIT DELHI
13 years ago

Dear anu,

Required period is pie..

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

Sagar Singh

B.Tech IIT Delhi



Lalitha bharadwaja
6 years ago
Period of mod(sinx)=pi=p1(say) and mod (cosx) = pi =p2(say)... Now as the given function is even function... Period(mod(sinx)+mod(cosx)=1/2*(LCM of p1,p2)= pi/2